OpenStudy (anonymous):

Simplify the following expression; express in terms of i. \[\sqrt{-72}\]

OpenStudy (lgbasallote):

\[i \sqrt{72}\] \[i \sqrt{(3)(3)(2)(2)(2)}\] \[(3)(2)(i)\sqrt{?}\] \[6i \sqrt{?}\]

OpenStudy (anonymous):

what ?